# A unified-description of curvature, torsion, and non-metricity of the metric-affine geometry with the Möbius representation

### v2

@@ Abstract
Keywords: gauge theories of gravity, metric-affine geometry, curvature, torsion, non-metricity, Möbius representation, Inönü-Wigner contraction
@@ 1 Introduction
General Relativity (GR), which is, on one hand, the most successful theory of gravity, employs the (pseudo-)Riemannian geometry to describe gravitational phenomena based on Einstein’s equivalence principle and the general covariance [1, 2]. On the other hand, Yang-Mills gauge field theory, which describes fundamental interactions in the elementary particle physics such as electromagnetic, weak, and strong interactions in a unified manner based on the gauge principle, makes use of a principal bundle with a specific gauge group for each theory [3, 4, 5, 6, 7, 8]. Inspired by these studies, Utiyama tried to establish a unified-description of all the above phenomena that is disciplined by the gauge invariant characteristics. [9] This is the dawn of the gauge theories of gravity. A detailed and historical development of this approach is illustrated in Refs [10, 11].
Metric-Affine gauge theories of Gravity (MAG) are one of such approach to gauge theories of gravity [12, 13, 14]. This theory is constructed upon the metric-affine geometry of which geometric quantities are not only curvature but also torsion and non-metricity (the latter one is further decomposed into dilation and shear). Therefore, MAG is an extension/modification of GR in terms of the gauge theories. In particular, as a modification of GR, a set of subclasses of MAG labeled as the Geometrical Trinity of Gravity (GTG) has been vigorously investigated by virtue of its equivalence in the equations of motion to GR up to boundary terms in the meaning of the action functional [15]. Under the imposition of the so-called Weitzenböch gauge condition [16, 17, 18, 19, 20], which implies that the curvature vanishes and then the geometry is flat, GTG has two subclasses: Symmetric Teleparallel Equivalent to GR (STEGR) and Teleparallel Equivalent to GR (TEGR).
@@ 2.1 Frame fields, internal space bundle, and Weitzenböch connection
This is the so-called Weitzenböch connection [16], which relates the connection 1-form to the affine connection via the (co-)frame fields and vice versa. Notice that Eq. (16) implies $\mathcal{D}_{\mu}e^{I}{}_{\nu}=0$, or equivalently, $\mathcal{D}_{\mu}e_{I}{}^{\nu}=0$. These properties are sometimes called vielbein postulate but, as shown in the above, always satisfied. That is, it is not a postulate.
@@ 6.1 Inönü-Wigner contraction
where $\epsilon^{(P)}$, $\epsilon^{(E)}$, $\epsilon^{(D)}$, and $\epsilon^{(S)}$ are a set of parameters and belong to the range $(0\,,1]$. Hereinafter, let us denote the parameter space spanned by the above parameters as $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ MA}=(\epsilon^{(P)}\,,\epsilon^{(E)}\,,\epsilon^{(D)}\,,\epsilon^{(S)})$, where “type of contraction/geometry in MA” denotes a geometry which is obtained by performing the contraction. Hereinafter, we use this notation. Then the above algebras are parameterized as follows:
@@ 6.3 An extension of metric-affine algebra and its contraction
The parameter space $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ EMA}=(0+\,,\epsilon^{(E)}\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ provides a set of new sort of geometric quantities. The coefficients of the third algebra of the second line and that of the second algebra of the third line in the algebra (69) diverge unless an appropriate set of the parameters $\epsilon^{(E)}$, $\epsilon^{(D)}$, and $\epsilon^{(S)}$ converges to zero simultaneously. The case of $\epsilon_{\rm\ abelian\ EMA}=(+0\,,+0\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ is the same situation as that of $\epsilon_{\rm\ abelian\ MA}=(+0\,,1\,,1)$ in the metric-affine case up to the difference of structure constants (local symmetry). Therefore, the intriguing cases are those of parameter spaces given by $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ EMA}=(+0\,,1\,,\epsilon^{(D)}\,,\epsilon^{(S)})$. If the contraction is performed by the parameters $\epsilon_{\rm\ EMA\ with\ shear-free}=(+0\,,1\,,1\,,+0)$ under the imposition of the gauge conditions (58) and (60) then the dilation given in Eq (70) remains but the shear given in Eq. (71) turns into ${\nearrow\!\!\!\!\!\!\!\Delta}=\omega^{(E)}\wedge\,\omega^{(S)}$. In the case of $\epsilon_{\rm\ EMA\ with\ dilation-free}=(+0\,,1\,,+0\,,1)$ under the same gauge conditions (58) and (60), the shear given in Eq. (71) remains but the dilation given in Eq. (70) turns into $\Delta=\omega^{(E)}\wedge\,\omega^{(D)}$. Remark, here, that the limitations of $\epsilon^{(D)}\rightarrow+0$ and $\epsilon^{(S)}\rightarrow+0$ are taken in the same order magnitude as $\epsilon^{(P)}\rightarrow+0$. The important point here is that these two algebras provide new geometric quantities, which are different from the dilation, the shear, or the non-metricity while holding the richness of the curvature and torsion 2-forms. This richness together with new geometric quantities would beyond the metric-affine geometry and yield a new class of gauge theories of gravity.
@@ 6.4 An extension of de Sitter/anti-de Sitter algebra and its contraction
The contraction of the above algebra generated by $\epsilon_{\rm\ Poincare\ in\ extended\ dS/AdS}=(+0\,,1\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ results in that of the algebra $\mathfrak{ema}(n\,,\mathbb{R})$ generated by $\epsilon_{\rm\ Poincare\ in\ EMA}=(\epsilon^{(P)}\,,1\,,\epsilon^{(D)}\,,\epsilon^{(S)})$, where $\epsilon^{(P)}$, $\epsilon^{(D)}$, and $\epsilon^{(S)}$ can be taken as an arbitrary value in $(0\,,1]$. So is for the contraction of algebras between that generated by $\epsilon_{\rm\ abelian\ Poincare\ in\ extended\ dS/AdS}=(+0\,,+0\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ in the extended dS/AdS algebra and that generated by $\epsilon_{\rm\ abelian\ Poincare\ in\ EMA}=(\epsilon^{(P)}\,,+0\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ in the extended metric-affine algebra under the choice of the structure constants as the Poincare algebra. Remark, here, that the contractions generated by $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ extended\ dS/AdS}=(1\,,+0\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ do not exist due to the divergence of the factor $(\epsilon^{(P)})^{2}/\epsilon^{(E)}$ of the first algebra in the first line. Therefore, it is enough to consider the case of $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ extended\ dS/AdS}=(1\,,1\,,\epsilon^{(D)}\,,\epsilon^{(S)})$.
@@ 7 Conclusions
As shown throughout the current paper, the fertility of the gauge theories of gravity should be ascribed to the structure of the internal space. The authors in Refs [29] unveiled that teleparallel gravity for the sake of cosmology, meaning that the geometry is homogeneous and isotropic, has five possible branches in the theory. This sort of classification play an important role in relating the number of propagating degrees of freedom revealed by the Dirac-Bergmann analysis to that number speculated by the cosmological perturbation. In fact, the authors in Ref [19] unveiled that the analysis of $f(T)$-gravity indicates the existence of five possible sectors and each sector generically has different propagating degrees of freedom. The relationships between the branches and the sectors are not unveiled today. Similarly, on one hand, the authors in Ref [61] unveiled that the analysis of coincident $f(Q)$-gravity indicates the existence of eight propagating degrees of freedom in a sector of the theory. On the other hand, the authors in Ref [47] showed that the theory has six propagating degrees of freedom in another sector of the theory, together with a possibility that the number is possibly seven or five in other sectors. From the viewpoint of cosmological perturbation, the authors in Ref [62] unveiled that seven propagating degrees of freedom exist in the non-trivial branch I. For the trivial and non-trivial branch II, the number could be four or less. The relationships between the branches and the sectors are also not unveiled today. In this regard, just scrutinizing curvature, torsion, and non-metricity themselves on the tangent bundle, which is likely to be a standard formalism in the present days, may be oversimplified. The fertility should be scrutinized in detail.

### v3

@@ Abstract
keywords:
gauge theories of gravity; metric-affine geometry; curvature; torsion; non-metricity; Möbius representation; Inönü-Wigner contraction.
{history}
@@ 1 Introduction
General Relativity (GR), which is, on one hand, the most successful theory of gravity, employs the (pseudo-)Riemannian geometry to describe gravitational phenomena based on Einstein’s equivalence principle and the general covariance [1, 2]. On the other hand, Yang-Mills gauge field theory, which describes fundamental interactions in the elementary particle physics such as electromagnetic, weak, and strong interactions in a unified manner based on the gauge principle, makes use of a principal bundle with a specific gauge group for each theory [3, 4, 5, 6, 7, 8]. Inspired by these studies, R. Utiyama tried to establish a unified-description of all the above phenomena that is disciplined by the gauge invariant characteristics [9]. This is the dawn of the gauge theories of gravity. A detailed and historical development of this approach is illustrated in Ref. [10].
Metric-Affine gauge theories of Gravity (MAG) are one of such approach to gauge theories of gravity [11, 12, 13]. This theory is constructed upon the metric-affine geometry of which geometric quantities are not only curvature but also torsion and non-metricity (the latter one is further decomposed into dilation and shear). Therefore, MAG is an extension/modification of GR in terms of the gauge theories. In particular, as a modification of GR, a set of subclasses of MAG labeled as the Geometrical Trinity of Gravity (GTG) has been vigorously investigated by virtue of its equivalence in the equations of motion to GR up to boundary terms in the meaning of the action functional [14]. Under the imposition of the so-called Weitzenböch gauge condition [15, 16, 17, 18, 19], which implies that the curvature vanishes and then the geometry is flat, GTG has two subclasses: Symmetric Teleparallel Equivalent to GR (STEGR) [20] and Teleparallel Equivalent to GR (TEGR) [21].
@@ 2.1 Frame fields, internal space bundle, and Weitzenböch connection
This is the so-called Weitzenböch connection [15], which relates the connection 1-form to the affine connection via the (co-)frame fields and vice versa. Notice that Eq. (16) implies $\mathcal{D}_{\mu}e^{I}{}_{\nu}=0$, or equivalently, $\mathcal{D}_{\mu}e_{I}{}^{\nu}=0$. These properties are sometimes called vielbein postulate but, as shown in the above, always satisfied. That is, it is not a postulate. Notice also that the covariant derivative $\nabla$ in Eq. (13) and $\mathcal{D}$ in Eq. (15) are equivalent, and, therefore, $\nabla$ satisfies the same rule as Eq. (12). Namely, a theory which is composed of using the covariant derivative $\nabla$ with respect to the Weitzenböch connection Eq. (16) is a gauge invariant theory under the (co-)frame transformation as discussed in Eq. (10).
@@ 6.1 Inönü-Wigner contraction
where $\epsilon^{(P)}$, $\epsilon^{(E)}$, $\epsilon^{(D)}$, and $\epsilon^{(S)}$ are a set of parameters and belong to the range $(0\,,1]$. Hereinafter, let us denote the parameter space spanned by the above parameters as
$\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ MA}=(\epsilon^{(P)}\,,\epsilon^{(E)}\,,\epsilon^{(D)}\,,\epsilon^{(S)})$, where “type of contraction/geometry in MA” denotes a geometry which is obtained by performing the contraction. Hereinafter, we use this notation. Then the above algebras are parameterized as follows:
@@ 6.3 An extension of metric-affine algebra and its contraction
The parameter space $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ EMA}=(0+\,,\epsilon^{(E)}\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ provides a set of new sort of geometric quantities. The coefficients of the third algebra of the second line and that of the second algebra of the third line in the algebra (69) diverge unless an appropriate set of the parameters $\epsilon^{(E)}$, $\epsilon^{(D)}$, and $\epsilon^{(S)}$ converges to zero simultaneously. The case of $\epsilon_{\rm\ abelian\ EMA}=(+0\,,+0\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ is the same situation as that of $\epsilon_{\rm\ abelian\ MA}=(+0\,,1\,,1)$ in the metric-affine case up to the difference of structure constants (local symmetry). Therefore, the intriguing cases are those of parameter spaces given by $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ EMA}=(+0\,,1\,,\epsilon^{(D)}\,,\epsilon^{(S)})$. If the contraction is performed by the parameters $\epsilon_{\rm\ EMA\ with\ shear-free}=(+0\,,1\,,1\,,+0)$ under the imposition of the gauge conditions (58) and (60) then the dilation given in Eq. (70) remains but the shear given in Eq. (71) turns into ${\nearrow\!\!\!\!\!\!\!\Delta}=\omega^{(E)}\wedge\,\omega^{(S)}$. In the case of $\epsilon_{\rm\ EMA\ with\ dilation-free}=(+0\,,1\,,+0\,,1)$ under the same gauge conditions (58) and (60), the shear given in Eq. (71) remains but the dilation given in Eq. (70) turns into $\Delta=\omega^{(E)}\wedge\,\omega^{(D)}$. Remark, here, that the limitations of $\epsilon^{(D)}\rightarrow+0$ and $\epsilon^{(S)}\rightarrow+0$ are taken in the same order magnitude as $\epsilon^{(P)}\rightarrow+0$. The important point here is that these two algebras provide new geometric quantities, which are different from the dilation, the shear, or the non-metricity while holding the richness of the curvature and torsion 2-forms. This richness together with new geometric quantities would go beyond the metric-affine geometry and yield a new class of gauge theories of gravity.
@@ 6.4 An extension of de Sitter/anti-de Sitter algebra and its contraction
The contraction of the above algebra generated by $\epsilon_{\rm\ Poincare\ in\ extended\ dS/AdS}=(+0\,,1\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ results in that of the algebra $\mathfrak{ema}(n\,,\mathbb{R})$ generated by $\epsilon_{\rm\ Poincare\ in\ EMA}=(\epsilon^{(P)}\,,1\,,\epsilon^{(D)}\,,\epsilon^{(S)})$, where $\epsilon^{(P)}$, $\epsilon^{(D)}$, and $\epsilon^{(S)}$ can be taken as an arbitrary value in $(0\,,1]$. So is for the contraction of algebras between that generated by $\epsilon_{\rm\ abelian\ Poincare\ in\ extended\ dS/AdS}=(+0\,,+0\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ in the extended dS/AdS algebra and that generated by ϵabelianPoincareinEMA=subscriptitalic-$\epsilon_{\rm\ abelian\ Poincare\ in\ EMA}=$
$(\epsilon^{(P)}\,,+0\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ in the extended metric-affine algebra under the choice of the structure constants as the Poincare algebra. Remark, here, that the contractions generated by $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ extended\ dS/AdS}=(1\,,+0\,,\epsilon^{(D)}\,,\epsilon^{(S)})$ do not exist due to the divergence of the factor $(\epsilon^{(P)})^{2}/\epsilon^{(E)}$ of the first algebra in the first line. Therefore, it is enough to consider the case of $\epsilon_{\rm\ type\ of\ contraction/geometry\ in\ extended\ dS/AdS}=(1\,,1\,,\epsilon^{(D)}\,,\epsilon^{(S)})$.
@@ 7 Conclusions
As shown throughout the current paper, the fertility of the gauge theories of gravity should be ascribed to the structure of the internal space. The authors in Ref [30] unveiled that teleparallel gravity for the sake of cosmology, meaning that the geometry is homogeneous and isotropic, has five possible branches in the theory. This sort of classification play an important role in relating the number of propagating degrees of freedom revealed by the Dirac-Bergmann analysis to that number speculated by the cosmological perturbation. In fact, the authors in Ref [18] unveiled that the analysis of $f(T)$-gravity indicates the existence of five possible sectors and each sector generically has different propagating degrees of freedom 333 In particular, Ref. [63] unveiled that $f(T)$-gravity has only one extra degrees of freedom in the comparision to the case of general relativity and, at a glance, it is inconsistent with the common knowledge that $f(T)$-gravity has five degrees of freedom in four-dimensional spacetime. In this point, Ref. [48] unveiled that the analysis in Ref. [63] is nothing but that in a generic sector which is also investigated in Ref. [18]. In detail, see Sec.V-A in Ref. [48]. . The relationships between the branches and the sectors are not unveiled today. Similarly, on one hand, the authors in Ref [64] unveiled that the analysis of coincident $f(Q)$-gravity indicates the existence of eight propagating degrees of freedom in a sector of the theory. On the other hand, the authors in Ref [48] showed that the theory has six propagating degrees of freedom in another sector of the theory, together with a possibility that the number is possibly seven or five in other sectors. From the viewpoint of cosmological perturbation, the authors in Ref [65] unveiled that seven propagating degrees of freedom exist in the non-trivial branch I. For the trivial and non-trivial branch II, the number could be four or less. The relationships between the branches and the sectors are also not unveiled today. In this regard, just scrutinizing curvature, torsion, and non-metricity themselves on the tangent bundle, which is likely to be a standard formalism in the present days, may be oversimplified. The fertility should be scrutinized in detail.
Finally, we briefly mention the possible extension of the method that is presented in the current paper from the viewpoints of $f(R)$-, $f(T)$-, and $f(Q)$-gravity. There are comprehensive reviews on each extended/modified theory of gravity: Refs. [66, 67], Ref. [68], and Refs. [69, 70], respectively. A common review on the generic framework of these theories is given in Ref. [71]. As mentioned in the previous paragraph, generically, the extended/modified theories of gravity have several sectors, and each sector has a proper constraint structure. This would lead to the proper physical phenomena for each sector. The crucial point for this bifurcation is symmetry breaking. GR has two fundamental symmetries: the local Lorentz symmetry, or generically speaking, the frame invariance, and the diffeomorphism symmetry. These symmetries are represented in terms of the Dirac analysis by the corresponding PB-algebra of the structure group $G$ of the internal bundle (See Sec. 2.1 and A) and the Hypersurface Deformation Algebra (HDA) [55] of the tangent bundle of a given spacetime manifold. $f(R)$-gravity holds these two sets of PB-algebras. This implies that the theory only has a single sector. In the case of $f(T)$-gravity, however, this theory loses the local Lorentz invariance, and this causes the bifurcation, although the HDA holds. Therefore, investigating further the pattern of the symmetry breaking, we would identify the relationships between the branches and the sectors mentioned in the previous paragraph. Then, we can perform the Inönü-Wigner contraction that corresponds to the breaking pattern of the PB-algebra and would find a comprehensive construction of the theory. It would be in the same situation as $f(Q)$-gravity. Furthermore, since the method presented in the current paper can extend the structure group freely to another one, it might be extended/modified to a theory beyond MAG. All of these investigations would be great future works.